# Asymmetrical Siamese Network for Point Clouds Normal Estimation

### v1

@@ Abstract
journal: Journal of  Templatesjournal: Journal of Computer-Aided Design
1 Introduction
@@
Currently, the mainstream normal estimation architectures [9, 10, 11, 12, 13, 14] based on deep learning are mainly inspired by PointNet [15] or the subsequent proposed network architectures [16, 17]. These architectures typically consist of three main parts: local patch alignment including PCA preprocessing or quaternion spatial transformer, feature extraction, and normal estimation including direct regression or fitting a local geometric surface. Table 1 provides more details about the comparisons between different architectures. Generally, except for PCA operations, effective feature representation extraction is necessary for each part of the architectures. For instance, in the quaternion spatial transformer operation within the first part, a global feature is extracted to estimate an accurate rigid rotation transformation for input space calibration. Similarly, in the normal estimation part, a global representation is also needed for direct normal estimation or point-wise weight prediction. Therefore, the performance of the networks significantly affects the accuracy and generalization ability of the normal estimation task. In this paper, we propose a novel Asymmetrical Siamese Network for normal estimation, which can be broadly applied on top of existing models such as DeepFit [18] and AdaFit [9]. Our Asymmetrical Siamese Network aims to align the multi-stage global features extracted from clean and noisy patches as closely as possible. This way, features from noisy patches can be enhanced by incorporating local geometric priors extracted from clean point clouds. Consequently, the learned features extracted by the networks can be more informative compared to the original models. By using prior information as guidance, we can improve the accuracy of normal estimation.
Table 1: A comparison of various normal estimation methods in the current mainstream literature.
@@ 2 Related Work
2.1 Conventional Normal Estimation
@@ 2.2 PCPNet and its Variants
2.3 Other Kinds of Methods
@@
2.4 Benchmarks for Normal Estimation
@@
3 Network Architecture
3.1 Overview
@@
3.2 Asymmetrical Siamese Network
@@
In addition, for a given patch, to ensure that the two asymmetric branches focus more on intrinsic noise information rather than differences in initial orientation, we apply a shared PCA operation in the preprocessing stage for the two sub-networks. Specifically, we first train the normal estimator $E_{1}$ (shown in the upper part of Fig. 1) for noise-free point cloud normal estimation. In this stage, for each noise-free patch, a corresponding noisy patch is required. The PCA operation is performed on the noisy patch to obtain an initial rotation transformation, which is then used for the initial canonical pose estimation of the noise-free patch in branch $E_{1}$. Once the first training process is completed, all weights of the normal estimator $E_{1}$ will be frozen. Subsequently, the second branch $E_{2}$ (the lower part in Fig. 1) is trained on noisy patches. The goal is to align the distribution of feature representations ($F_{1}^{1}$, $F_{2}^{1}$, $F_{1}^{2}$ and $F_{2}^{2}$ denoted in Fig. 1) extracted from both branches by optimizing the feature matching distance. Finally, in the inference stage, the second branch $E_{2}$ can perform the normal estimation task independently. Experiments in Sec. 5.2 demonstrate that the feature representations obtained using the proposed feature matching strategy are more effective than the global features directly extracted from a single branch.
@@ 3.3 Loss Function
Table 2: Comparison of the RMSE angle error for unoriented normal estimation of our Asymmetrical Siamese Architecture to the classical deep learning methods on datasets PCPNet Dataset. Here, “w” indicates using our method, while “w/o” refers to the original methods.
@@
Table 3: Comparison of the RMSE angle error for unoriented normal estimation of our Asymmetrical Siamese Architecture to the classical deep learning methods on datasets FamousShape Dataset. Here, “w” indicates using our method, while “w/o” refers to the original methods.
@@
4 Multi-View Normal Estimation Dataset
@@
5 Experiments
5.1 Datasets and Settings
@@
Table 4: Comparison of RMSE angle error for normal estimation training, with and without the incorporation of our multi-view dataset, on our detailed models test set from the Multi-View Normal Estimation Dataset. Here, “w” indicates the utilization of both our multi-view and PCPNet training datasets, whereas “w/o” denotes training solely with the PCPNet training dataset.
@@
5.2 Enhancements in Asymmetrical Siamese Architecture
To validate the effectiveness of our Asymmetrical Siamese framework, we compared it with several mainstream normal estimation algorithms, including DeepFit [12], AdaFit [9], and GraphFit [14]. To ensure the fairness of the comparison experiments, we used the same parameter settings as in the original works, with the only difference being that we trained two branches using these methods as the backbone. Specifically, we used the noise-free data from the PCPNet training set for the first training stage and the complete PCPNet training set for the second stage of the Asymmetrical Siamese Network. We assessed the accuracy improvements on the PCPNet and FamousShape datasets with and without the Asymmetrical Siamese framework through both quantitative and qualitative comparisons. As shown in Tab. 2, our method brings improvements to the DeepFit and AdaFit models on the PCPNet dataset. Although there is a slight decrease in performance for the GraphFit method, our approach consistently enhances performance for data with high noise level across all methods. Then, Tab. 3 demonstrates that, after training with the Asymmetrical Siamese framework, our method also improves performance on the FamousShape dataset, which is different from the PCPNet dataset. This indicates that our training approach with two stages can indeed enhance the generalization ability of the models by conveying pattern information from noise-free data. Additionally, the AUC results are illustrated in Fig. 3, where our method consistently demonstrates superior performance, indicating its remarkable stability across different angular thresholds. The enlarged part of Fig. 3 shows that training with our Asymmetric Siamese Network brings noticeable improvements, particularly as the error increases.
@@ 5.2 Enhancements in Asymmetrical Siamese Architecture
Table 5: Comparison of RMSE angle error for normal estimation training on our CAD test set from the Multi-View Normal Estimation Dataset, with and without the incorporation of our multi-view dataset. “w” indicates the utilization of both our multi-view and PCPNet training datasets, whereas “w/o” denotes training solely with the PCPNet training dataset.
@@
Table 6: Comparison of RMSE angle error for normal estimation training, with and without the incorporation of our multi-view dataset, on the PCPNet test dataset. Here, “w” indicates the utilization of both our multi-view and PCPNet training datasets, whereas “w/o” denotes training solely with the PCPNet training dataset.
@@
5.3 Domain Gap Between Datasets
In this section of the study, we explored the impact of differences between datasets on model performance. Firstly, we evaluated the multi-view test dataset using existing mainstream methods. As shown in Tab. 5 and Tab. 4 , results from testing solely with the PCPNet dataset (“w/o”) indicate poorer performance on our proposed dataset, including CAD models and detailed models. This highlights significant differences between the two datasets. However, after retraining existing models with a portion of our proposed dataset, significant improvements are observed on the multi-view scanning test dataset proposed in this paper, as indicated in the “w” column of Tab. 5 and Tab. 4. In addition, Fig. 6 and Fig. 7 also provide visual comparisons of the performance enhancement on CAD models and detailed models after training with the new dataset.
Additionally, we compared the performance of models trained with multi-view point cloud data against the original models on the PCPNet dataset, as shown in Tab. 6. The experimental results indicate that incorporating multi-view data for training significantly improves the performance of the PCPNet and DeepFit models. However, for models with stronger learning capabilities, training with additional data that has a significant domain gap results in a slight performance decline on the PCPNet dataset. This suggests an unsettling fact: the high-performing normal estimation models including AdaFit [9], GraphFit [14] and Hsurf [13] may actually be in an overfitted state.
Table 7: Comparisons of RMSE angle error on the asymmetric Siamese DeepFit with different feature matching weight values.
@@ 5.3 Domain Gap Between Datasets
5.4 More Analysis for Asymmetrical Siamese Architecture
Parameter Setting. Considering that the feature matching weights used in our Asymmetric Siamese Architecture will be utilized to balance the results of normal estimation with the matching degree of noise-free features, which affects the model’s generalization ability, the choice of weight values should be carefully discussed and analyzed. In our framework, the global features from the QSTN rotation process and the final latent code are involved in the matching constraints. Therefore, we compared the impact of these two matching losses on the model’s accuracy. As shown in Tab. 7, our experiments use the DeepFit [12] model as the backbone, with the matching learned features in the QSTN module belonging to the $l_{1}$ layer and the final patch’s features to the $l_{2}$ layer. The experimental results indicate that the optimal matching weight for the QSTN layer is 0.5. On the other hand, matching the global features of the patch degrades the model’s performance. This highlights the critical importance of consistency in the QSTN rotation process.
@@ 6 Conclusion
Acknowledgment
We would like to thank all the reviewers for their valuable comments and feedback. This work was supported by NSFC (No. 62002040) and Fundamental Research Funds for the Central Universities (No. 3132024271)

### v2

@@ Abstract
I Introduction
@@
Currently, the mainstream normal estimation architectures [9, 10, 11, 12, 13, 14] based on deep learning are mainly inspired by PointNet [15] or the subsequent proposed network architectures [16, 17]. These architectures typically consist of three main parts: local patch alignment including PCA preprocessing or quaternion spatial transformer, feature extraction, and normal estimation including direct regression or fitting a local geometric surface. Tab. I provides more details about the comparisons between different architectures. Generally, except for PCA operations, effective feature representation extraction is necessary for each part of the architectures. For instance, in the quaternion spatial transformer operation within the first part, a global feature is extracted to estimate an accurate rigid rotation transformation for input space calibration. Similarly, in the normal estimation part, a global representation is also needed for direct normal estimation or point-wise weight prediction. Therefore, the performance of the networks significantly affects the accuracy and generalization ability of the normal estimation task. In this paper, we propose a novel Asymmetrical Siamese Network for normal estimation, which can be broadly applied on top of existing models such as DeepFit [18] and AdaFit [9]. Our Asymmetrical Siamese Network aims to align the multi-stage global features extracted from clean and noisy patches as closely as possible. This way, features from noisy patches can be enhanced by incorporating local geometric priors extracted from clean point clouds. Consequently, the learned features extracted by the networks can be more informative compared to the original models. By using prior information as guidance, we can improve the accuracy of normal estimation.
TABLE I: A comparison of various normal estimation methods in the current mainstream literature.
@@ II Related Work
II-A Conventional Normal Estimation
@@ II-B PCPNet and its Variants
II-C Other Kinds of Methods
@@
II-D Benchmarks for Normal Estimation
@@
III Network Architecture
III-A Overview
@@
III-B Asymmetrical Siamese Network
@@
In addition, for a given patch, to ensure that the two asymmetric branches focus more on intrinsic noise information rather than differences in initial orientation, we apply a shared PCA operation in the preprocessing stage for the two sub-networks. Specifically, we first train the normal estimator $E_{1}$ (shown in the upper part of Fig. 1) for noise-free point cloud normal estimation. In this stage, for each noise-free patch, a corresponding noisy patch is required. The PCA operation is performed on the noisy patch to obtain an initial rotation transformation, which is then used for the initial canonical pose estimation of the noise-free patch in branch $E_{1}$. Once the first training process is completed, all weights of the normal estimator $E_{1}$ will be frozen. Subsequently, the second branch $E_{2}$ (the lower part in Fig. 1) is trained on noisy patches. The goal is to align the distribution of feature representations ($F_{1}^{1}$, $F_{2}^{1}$, $F_{1}^{2}$ and $F_{2}^{2}$ denoted in Fig. 1) extracted from both branches by optimizing the feature matching distance. Finally, in the inference stage, the second branch $E_{2}$ can perform the normal estimation task independently. Experiments in Sec. V-B demonstrate that the feature representations obtained using the proposed feature matching strategy are more effective than the global features directly extracted from a single branch.
@@ III-C Loss Function
TABLE II: Comparison of the RMSE angle error for unoriented normal estimation of our Asymmetrical Siamese Architecture to the classical deep learning methods on datasets PCPNet Dataset. Here, “w” indicates using our method, while “w/o” refers to the original methods.
@@
TABLE III: Comparison of the RMSE angle error for unoriented normal estimation of our Asymmetrical Siamese Architecture to the classical deep learning methods on datasets FamousShape Dataset. Here, “w” indicates using our method, while “w/o” refers to the original methods.
@@
IV Multi-View Normal Estimation Dataset
@@
V Experiments
V-A Datasets and Settings
@@
TABLE IV: Comparison of RMSE angle error for normal estimation training, with and without the incorporation of our multi-view dataset, on our detailed models test set from the Multi-View Normal Estimation Dataset. Here, “w” indicates the utilization of both our multi-view and PCPNet training datasets, whereas “w/o” denotes training solely with the PCPNet training dataset.
@@
V-B Enhancements in Asymmetrical Siamese Architecture
To validate the effectiveness of our Asymmetrical Siamese framework, we compared it with several mainstream normal estimation algorithms, including DeepFit [12], AdaFit [9], and GraphFit [14]. To ensure the fairness of the comparison experiments, we used the same parameter settings as in the original works, with the only difference being that we trained two branches using these methods as the backbone. Specifically, we used the noise-free data from the PCPNet training set for the first training stage and the complete PCPNet training set for the second stage of the Asymmetrical Siamese Network. We assessed the accuracy improvements on the PCPNet and FamousShape datasets with and without the Asymmetrical Siamese framework through both quantitative and qualitative comparisons. As shown in Tab. II, our method brings improvements to the DeepFit and AdaFit models on the PCPNet dataset. Although there is a slight decrease in performance for the GraphFit method, our approach consistently enhances performance for data with high noise level across all methods. Then, Tab. III demonstrates that, after training with the Asymmetrical Siamese framework, our method also improves performance on the FamousShape dataset, which is different from the PCPNet dataset. This indicates that our training approach with two stages can indeed enhance the generalization ability of the models by conveying pattern information from noise-free data. Additionally, the AUC results are illustrated in Fig. 3, where our method consistently demonstrates superior performance, indicating its remarkable stability across different angular thresholds. The enlarged part of Fig. 3 shows that training with our Asymmetric Siamese Network brings noticeable improvements, particularly as the error increases.
@@ V-B Enhancements in Asymmetrical Siamese Architecture
TABLE V: Comparison of RMSE angle error for normal estimation training on our CAD test set from the Multi-View Normal Estimation Dataset, with and without the incorporation of our multi-view dataset. “w” indicates the utilization of both our multi-view and PCPNet training datasets, whereas “w/o” denotes training solely with the PCPNet training dataset.
@@
TABLE VI: Comparison of RMSE angle error for normal estimation training, with and without the incorporation of our multi-view dataset, on the PCPNet test dataset. Here, “w” indicates the utilization of both our multi-view and PCPNet training datasets, whereas “w/o” denotes training solely with the PCPNet training dataset.
@@
V-C Domain Gap Between Datasets
In this section of the study, we explored the impact of differences between datasets on model performance. Firstly, we evaluated the multi-view test dataset using existing mainstream methods. As shown in Tab. V and Tab. IV , results from testing solely with the PCPNet dataset (“w/o”) indicate poorer performance on our proposed dataset, including CAD models and detailed models. This highlights significant differences between the two datasets. However, after retraining existing models with a portion of our proposed dataset, significant improvements are observed on the multi-view scanning test dataset proposed in this paper, as indicated in the “w” column of Tab. V and Tab. IV. In addition, Fig. 6 and Fig. 7 also provide visual comparisons of the performance enhancement on CAD models and detailed models after training with the new dataset.
Additionally, we compared the performance of models trained with multi-view point cloud data against the original models on the PCPNet dataset, as shown in Tab. VI. The experimental results indicate that incorporating multi-view data for training significantly improves the performance of the PCPNet and DeepFit models. However, for models with stronger learning capabilities, training with additional data that has a significant domain gap results in a slight performance decline on the PCPNet dataset. This suggests an unsettling fact: the high-performing normal estimation models including AdaFit [9], GraphFit [14] and Hsurf [13] may actually be in an overfitted state.
TABLE VII: Comparisons of RMSE angle error on the asymmetric Siamese DeepFit with different feature matching weight values.
@@ V-C Domain Gap Between Datasets
V-D More Analysis for Asymmetrical Siamese Architecture
Parameter Setting. Considering that the feature matching weights used in our Asymmetric Siamese Architecture will be utilized to balance the results of normal estimation with the matching degree of noise-free features, which affects the model’s generalization ability, the choice of weight values should be carefully discussed and analyzed. In our framework, the global features from the QSTN rotation process and the final latent code are involved in the matching constraints. Therefore, we compared the impact of these two matching losses on the model’s accuracy. As shown in Tab. VII, our experiments use the DeepFit [12] model as the backbone, with the matching learned features in the QSTN module belonging to the $l_{1}$ layer and the final patch’s features to the $l_{2}$ layer. The experimental results indicate that the optimal matching weight for the QSTN layer is 0.5. On the other hand, matching the global features of the patch degrades the model’s performance. This highlights the critical importance of consistency in the QSTN rotation process.
@@ VI Conclusion
Biography
[Uncaptioned image] Wei Jin received her B.S. and M.D. degrees in computational mathematics from Dalian University of Technology, China, in 2012 and 2015, respectively. She is currently a lecturer at Dalian Neusoft University of Information, China. Her research interests include computer graphics, image processing, and machine learning.
[Uncaptioned image] Jun Zhou received his B.S. and Ph.D. degrees in computational mathematics from Dalian University of Technology, China, in 2013 and 2020, respectively. He is currently a lecturer in the College of Information Science and Technology at Dalian Maritime University, China. His research interests include computer graphics, image processing, and machine learning.
[Uncaptioned image] Nannan Li is an associate professor in School of Information Science and Technology at Dalian Maritime University. She received her B.S. and M.D-Ph.D in Computational Mathematics at Dalian University of Technology. Her research interests include computer graphics, differential geometry analysis and processing, computer aided geometric design and machine learning.
[Uncaptioned image] Haba Madeline holds a Bachelor’s Degree in Computer Science and Technology from Shenyang University of Technology, awarded in 2022. She is currently pursuing a Master’s Degree in Software Engineering at Dalian Maritime University. Her research interests include Computer Vision and Pattern Recognition.
[Uncaptioned image] Xiuping Liu received the BSc degree from Jilin University, China, in 1990, and the PhD degree from the Dalian University of Technology, China, in 1999, respectively. She is a professor with the Dalian University of Technology. Between 1999 and 2001, she conducted research as a postdoctoral scholar in the School of Mathematics, Sun Yat-sen University, China. Her research interests include shape modeling and analyzing.